A firm plans to make a change that will increase sales from 22,000 vases to 27,000 vases and increase its average collection period from 25 days to 38 days. The price per vase is $400 and the variable cost per unit is $325. The required return on similar-risk investments is 15%.

1.What is the additional profit contribution from sales if the change is made?

2.What is the cost of the marginal investment in accounts receivable for this change?

3.What is the net profit from this change?

4.Should the firm make the change? Explain.
Answer #1

Detailed solution is shown below ask if any doubt in comment

. Additional profit contribution = (27000-2200) (400-325) = 5000 (75) Additional poofit contribution = $375000 Average AccounI net profit - Additional cost of profit Margonal investment 375000 - 63575 net profit = $311425 h Yes, The fiors should make